You take a frisbee to the top of the Washington Monument and send it sailing along horizontally at a speed of 5 m/s from the very top of the structure, which is 169.2 m above the ground.

A) How long will the frisbee remain in the air? (Remember--perfect physics world--no air resistance!)

 

B) How far down range will the frisbee go (down range - horizontal distance.)?

 


1
Expert's answer
2021-05-11T17:21:20-0400

(A) We can find the time during which the frisbee will remain in the air from the kinematic equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}},""t=\sqrt{\dfrac{2\cdot169.2\ m}{9.8\ \dfrac{m}{s^2}}}=5.87\ s."


(B) We can find the range of the frisbee from the kinematic equation:


"x=v_0t=5\ \dfrac{m}{s}\cdot5.87\ s=29.35\ m."